OpenStudy (anonymous):

2x-6y=30 3x-9y=-27 Solve the system of equations by graphing ( i don't need the graphing done, I just need a walk through of how to solve this)

OpenStudy (anonymous):

okay to do this first that the 1st eq'n 2x -6y = 30 and graph it. I would graph this line by finding the x and y intercepts x int (let y = 0): (15,0) y int (lets x = 0): (0,-5) next, take the second equation 3x-9y=-27 and graph it as well x int (let y = 0): (-9,0) y int (lets x = 0): (0,3) wherever the lines meet is the sol'n

OpenStudy (anonymous):

How did you come up with the 15 and -5??

OpenStudy (anonymous):

for 15, the x intercept occurs when y = 0. Sub y = 0 into the equation -2x -6(0) = 30 -2x = 30 x = -25 the y intercept occurs when x = 0. Sub x = 0 into the equation -2(0) -6y = 30 -6y = 30 y = -5

OpenStudy (anonymous):

ooohhh I see...okay, so since there is more than one ordered pair, that means there is more than one solution?

OpenStudy (anonymous):

The solution will be one point on the graph (x,y) where the two lines intercept. The lines should only cross once. The (15,0) and (0,-5) are just two points so you can graph the first line.

OpenStudy (anonymous):

OOOHHHHHHH DUH!!! lol

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

my graph should look like an X ?

OpenStudy (anonymous):

would the ordered pair be (0,3)

OpenStudy (anonymous):

when these two equations are graphed, they are parallel; which means no solution.